The sum of the ages of a son and his father is 35 and the product of their ages is 150. Find their ages.

Answer by Cromlix(4381)

The sum of the ages of a son and his father is 35
x + y = 35
the product of their ages is 150.
xy = 150
x = 150/y
Substitute x = 150/y into
x + y = 35
150/y + y = 35
Common denominator = y
150 + y^2 = 35y
y^2 - 35y + 150 = 0
(y - 5)(y - 30)
Son's age is 5 years old.
The father's age is 30 years old.
Hope this helps.
:-)